# JEL Classification: C51, C81, D83, O13, Q12, Q54

### v1

@@ 3.1.3 Creation of District-Level EO Data Set
The creation of the district-level panel data follows the workflow summarized in FigureΒ 3. We start by using the rice area maps for Aman (wet) season for each year to mask the flood maps, so that we only consider inundation levels for rice area. We also use the rice maps to mask EVI data from MODIS so that we only consider EVI values for rice area.
@@
Given the resolution of the data it is not possible to measure EVI on a rice field for use in predicting yields as in Lobell etΒ al. (2020). The rice maps exist at a spatial resolution of 250m while the mean size of a rice plot in the RMS is 0.68 ha. Nor is it possible to measure the exact, agronomically relevant inundation of a rice field, given that the flood maps exist at 500m resolution. To overcome this data limitation, we calculate a variety of EVI and flood metrics to ensure our proxies for relevant yield and flooding are robust. For EVI, we calculate the maximum value of EVI for the season, the minimum, the mean and median values. All of these metrics turn out to by highly correlated with each other, with correlation coefficients ranging from 0.75 up to 0.99, depending on the specific comparison. Because of this, all of the analysis in the paper is based on our preferred metric, the maximum value of EVI. Results are strongly robust to use of any of the other metrics.
For flood, we calculate four metrics. First, 1) cumulative flooding, which is measured as the sum of the weekly fractional pixel values for the growing season (June to December) for a given year. We also calculate 2) the maximum fractional values for the season as well as the 3) mean and 4) median values. Unlike with the EVI metrics, there is more variation in the flood data meaning the flood metrics are not as highly correlated (though mean and median are highly correlated). As one will see, our results are generally robust to the use of flood measure, though this is not always the case. Because of this, throughout the paper we present results using all four flood metrics.
@@
Having derived our measures of EVI and flooding from the remote sensing time series, we when take the average value of that metric for each pixel within a district. This provides us with district-level values for each of our EO based metrics. We then combine this with the administrative data on seed dissemination. From IRRI and the Bangladesh Rice Research Institute (BRRI), we developed an exhaustive list of all seed production organizations, both private and public. While both IRRI and BRRI have been instrumental in developing submergence tolerant seeds varieties, neither is engaged in seed multiplication, which is conducted by a number of public and private entities. Investment in the seed system helped promote and disseminate STRVs from 2010 through 2019. We visited each seed producer/distributor and obtained disaggregate (district) level data on 1) the STRV variety names being multiplied or sold/distributed in the district, 2) the amount (tons) of each variety being multiplied in the district, and 3) the amount (tons) of each variety being sold/distributed in the district.999Private organizations sell seed while public organizations may sell the seed or distribute for free. The administrative data provides a proxy for the availability of STRV seeds in each district in each year starting in 2010. A limitation of this data is that farmers may recycle seed or obtain seed from outside their district. We use the cumulative amount of seed available in each district in each year to try and account for the growing availability of STRVs.
3.2 Household Survey Data
To investigate the impact of adoption of STRVs on household welfare, we use data from the Rice Monitoring Survey (RMS). The RMS is a BMGF project designed to capture varietal turnovers over time in Bangladesh, India, and Nepal. The data was originally collected as two waves of a panel in 2014 and 2017 (Yamano, 2017). Households were selected following a clustered random sampling procedure to ensure the overall survey was representative of rice growing regions of each country. The plan was to collect a third wave in 2020, but data collection efforts were delayed due to COVID until 2022 and only Bangladesh was revisited. In total, 1,500 households were part of the initial sampling frame in Bangladesh. The RMS was able to follow-up with 1,490 households in 2017 and 1,484 households in 2022. This gives us an attrition rate of over a span of eight years.
We construct an unbalanced panel of households that cultivated rice in Aman season. This gives us a total of 3,865 observations from 1,488 households. Our primary outcome of interest is rice yield at the household level, which we construct by summing up rice harvest (kg) on all plots and dividing by the sum of area (ha) for all rice plot. We consider a household as having adopted STRVs if they report planting any flood tolerant rice variety on at least one plot. To measure the incidence of flooding faced by the household, we combine household GPS location with the EO flood maps we generate.101010While we would ideally use plot locations to match with flood data, only household GPS locations were collected in 2014 and 2017. Thus, we only know the precise location of plots in 2022. We then calculate five measures of flooding for each survey year: cumulative, maximum, mean, median, and a measure of the flood experienced by the village. Our village measure includes cumulative flooding for all households in the village, minus the household in question. We also construct flood time series for each household going back to 2002. In some specifications we use the householdβs historic experience with flooding to calculate the probability of a household experiencing a flood in a given year as an instrument for STRV adoption.
@@ 3.3 Data Limitations
Before moving on to our empirical method, it is important to note and discuss the data limitations in the study. In an ideal situation, researchers would have collected baseline data from rice growing households prior to the release of STRV seeds and would have followed-up with these households periodically over the years. Unfortunately, no such panel data with pre-release baseline exists. This motivates our use of EO data to reconstruct how the EVI signal responds to flooding pre-STRV release and then to examine how that signal changes post release.
@@
In terms of ideal EO data, one would like data from a sensor that is at both a high spatial and temporal resolution and that can penetrate clouds. Additionally, one would like to have ground truth data in terms of both flooding and rice area/yields across both space and time. As highlighted in SectionΒ 3.1, such data does not exist, mainly because high resolution, cloud penetrating EO sensors did not exist in the public prior to the introduction of STRVs. Thus, there is no way to construct a baseline at both a high spatial and temporal resolution. Additionally, baseline ground truth data do not exist because no baseline survey was conducted. This poverty of ideal data motivates our approach to constructing rice area maps using Google Earth images and flood maps using a fusion model.
@@
In developing our approach, we prioritized products with a high temporal resolution (MODIS) at the expense of products with a high spatial resolution (Landsat) because of the importance of capturing transitory flood events and EVI values immediately prior to harvest. We are aware of this trade-off and the limitation it creates: our EO data set is constructed at 500m, which is an area equal to 25 ha. In a cropping system like rice in Bangladesh, where the average rice area for a household is less than one ha, this spatial resolution is not ideal. However, considering the types of measurement error introduced by the trade-off leads us to believe that prioritizing temporal over spatial resolution is the preferred approach. We have explored conducting the analysis using Landsat 30m data. However, Landsatβs temporal resolution is 16 days, meaning it frequently misses known flood events and EVI readings at harvest are frequently missed. Given this temporal resolution, it is possible to under count by missing a flood or harvest but it is not possible to over count. This means that the measurement error introduced is non-classical and while the direction of the bias is known, the size of the bias is unknown. By contrast, the relatively course spatial resolution of MODIS, combined with the ground truth data, and our machine learning classification approach means that we may misallocate a pixel to rice or no-rice and to flood or no-flood, but that misallocation is likely random. This means the measurement error introduced should be classical - increasing the variance or noise in the data but not introducing any systematic bias. Essentially, the spatial fuzziness of MODIS allows us to make guided inference on whether or not a pixel is rice/non-rice or flooded/not-flooded while the missing temporal images of Landsat do not yield to making inference about what happened in those missing pixels. This spatial fuzziness can be measured in terms of accuracy metrics for our classification algorithms (RF and CNN-LSTM), which are in the range of - (see Online AppendixΒ C for more details).
In terms of data limitation to the RMS, as adoption of STRVs happen on a plot-by-plot and crop-by-crop basis, the ideal data would be plot-level panel data. Additionally, one would want historic and contemporaneous flood data at a high enough spatial and temporal resolution to measure flooding on that specific plot and not a neighboring plot. Unfortunately, in the RMS, it is not possible to track plots over time. Plot dimensions shift from year-to-year and farmers buy, sell, and rent plots from season-to-season. Additionally, GPS coordinates for plots were not collected as part of the RMS in 2014 and 2017, so we do not know the exact locations of plots in those years. Even if we did, like we do in 2022, the EO flood data is not at a high enough spatial resolution to measure flooding on one specific plot. We do, as a robustness check, conducted our analysis on the pooled plot data and compare these results to those from the household-panel.
@@
Finally, we also lack objectively measured harvest information in the EO data and both harvest and rice area in the RMS data. In the EO data, we use EVI as a proxy for harvest, though this approach is not without its critics. Much of the research demonstrates strong correlation between EVI, and particularly peak EVI, and yields (Mkhabela etΒ al., 2011; Liu etΒ al., 2020; Lobell etΒ al., 2020). However, the accuracy of using EVI as a proxy for yields varies substantially by location (Azzari etΒ al., 2017) and crop, with EVI performing particularly poorly for rice in the U.S. (Johnson, 2016). For rice in Bangladesh, the correlation between EO derived metrics and yield can vary from - (Islam etΒ al., 2021). However, Islam etΒ al. (2021) relied on government reported area yield statistics, a notoriously inaccurate source for accurate and objective measures of yield in low-income countries (Kilic etΒ al., 2021), where statistical departments are often under-resourced and politicized. We are unaware of tests of accuracy of EVI with rice yields in Bangladesh that rely on objective measures of yield.
In the household survey data, we are forced to rely on self-reported measures of rice area and harvest. Since Carletto etΒ al. (2015), it has become standard to use handheld GPS devices to calculate land area for at least some plots. And since Abay etΒ al. (2019), it has become common to conduct crop-cuts on at least some plots. While we did conduct GPS measures of plots in the 2022 round, these methods were not common when the first two rounds of RMS data were collected. As Abay etΒ al. (2019) shows that correcting for biased self-reporting in one (land area or harvest weight) but not the other can actually aggravate the bias, and as we have objectively measured land area for one year and lack objectively measured harvest in all years, we have decided to rely solely on the self-reported values in all years.
@@
While these data limitations are not unique to our study, and in fact would be present in any attempt to measure impact of an agricultural technology introduced over a decade ago, they should be kept in mind when interpreting our results.
4 Research Design
4.1 Pre-Analysis Plan
In developing the following empirical approach, we pre-specified our analysis and registered it with Open Science Framework (OSF) (AlΒ Rafi etΒ al., 2023).111111Note that the PAP was registered on 9 February 2023. While in this paper we refer to the third wave of panel data as 2022, this is because we refer to all cropping season data by the year of planting not the year of harvest. Thus data from 2014, 2017, and 2022 all come from the Aman season in which planting occurred in 2014, 2017, or 2022 and harvest at the end of that calendar year, extending into the subsequent calendar year. Data collection for 2022 was not completed until 12 March 2023. Timestamps on the data are available for verification. While pre-analysis plans (PAPs) are common in lab and field experiments, they remain relatively uncommon for observational studies. However, Josephson and Michler (2023) argue that any studying collecting new data, regardless of the type of data, or trying to establish causal effects should develop and register a PAP. They point out that the very first PAP in economics was observational in nature and was created by Neumark (2001), who was studying the impacts of changes to minimum wage on employment - a particularly contentious issue - and one in which a PAP could be developed ahead of an announced wage increase.
Our study, though observational, is well suited to a PAP given that the research relies on a newly collected round of data and that it uses DID and IV methods to establish causal effects of STRV adoption. In addition to preventing the -hacking that Brodeur etΒ al. (2020) shows can be common to DID and IV methods, the plan also helps limit research degrees of freedom - the ability of researchers to make many reasonable or justifiable decision during the data cleaning and analysis process. In our context, it is unclear what the best EO proxy is for agronomically relevant flooding or for yield. It is also unclear what the best IV for adoption might be. There are also a large number of reasonable criteria to use for deciding how to deal with outliers, how to convert, scale, and aggregate data, and how to deal with missing values. In the absence of a PAP, a researcher might be tempted to try many alternative options in these domains and then only present the choices that result it strongly significant results. Our PAP ties our hands so that we can avoid the temptation to selective report results favorable to our hypothesis.
Throughout the results section, where data definitions, approaches, or inference criteria differ from our plan, we highlight these differences. The most substantial deviation from the PAP is that we initially planned to measure flooding as the total number of distinct flood events. Since we filed our PAP, the research team developed the fusion model described in Giezendanner etΒ al. (2023) that provides a fractional index of inundation in a pixel. This new approach provides much greater opportunity to develop more precise measures of flooding than the original binary indicator of flooding on a given day. Following advice in Duflo etΒ al. (2020) and Janzen and Michler (2021) to not stick to an inferior pre-specified method when a new, superior method becomes available, we adopted the fractional flood index data instead of the binary data. To help alleviate concerns that we selectively report a new flood measure that is particularly favorable to our hypotheses, we report results using four different flood measures based on the new data.
4.2 Empirical Method
Our empirical method follows that laid out in Gollin etΒ al. (2021) to estimate the impact of Green Revolution technologies on yields. We first estimate the effects of the availability of STRV seeds at the district-level on relative rice yields as proxied by EVI with and without flooding. Our empirical strategy relies on the release dates of various varieties and the staggered accumulation of seed in each district, which we argue is plausibly exogenous to rice yields. We then estimate the effects of adoption of STRV seed by a household on relative rice yields with and without flooding. We take a variety of approaches to address the endogeneity of the adoption decision. These include using a fixed effects with-in estimator as well as instrumental variables based on the probability that a households would experience a flood. We enrich these results by also estimating the impact of adoption using the RMS cross-sectional plot-level data.
4.2.1 EO District-Level Framework
We estimate the effects of STRVs on EVI as a proxy for yields using annual data at the district-level. We start be estimating a simple event study model:
@@ 4.2.1 EO District-Level Framework
where indexes districts and indexes time. The two terms and denote district and time fixed effects, meaning that only within-district time variation in relative EVI remains. The district fixed effects control for all district-specific time-invariant variation and the time fixed effects control for all time-specific district-invariant variation. We also include controls for the percentage of a district that is under rice cultivation and the cumulative flooding in the district, both of which vary by district and time.
@@
There are two shortcomings to the event study approach. One is that STRVs reach different districts in different years. It is plausible that STRV seeds reached more flood-prone districts, or districts with more advanced farming techniques, earlier than other districts. If these characteristics are time-invariant, our district fixed effects will pick them up. However, one can imagine situations where time-variant, district-specific events accelerated the arrival of STRV seed. Such a situation would be if public or private seed multipliers rushed STRV seed to a district in a year immediately following an unexpected flood event. While we explicitly control for contemporaneous flooding, flood events in the recent past might make the year a district got STRVs no longer exogenous. A second shortcoming is that STRVs are likely to increase EVI only very slightly over time, as a change in EVI would only occur when there was a flood lasting more than seven but no more than 17 days. If a district did not experience substantial flooding in a year after STRVs were released, we would not expect any change in EVI. Thus, the event study will under estimate the effects of STRV release date on EVI as STRVs will only effect EVI if there was a flood.
@@
To address these two issue, we next estimate a simple DID model:
@@
where is an indicator equal to one in the years after the first introduction of STRVs in 2010 and zero otherwise. Similarly, equals one in districts that are prone to flooding and zero in districts that are not prone to flooding. To define a district as flood-prone or not, we first calculate the median amount of flooding across all districts and all years. We then calculate the median flood value in each district across all years. We categorize a district as flood-prone if its median value is above the median value in all the data. The remaining terms are as defined in EquationΒ (4.1).121212Note that because we have converted both adoption and flood to binary indicators that do not vary over time, our DID estimator is not subject to the recently exposed potential biases of DID in complicated settings. For a summary of what settings complicate DID estimation and the numerous ways to correct for bias, see Roth etΒ al. (2023).
@@
We expect EVI in flood-prone districts to increase after 2010, which was when STRVs first became available in Bangladesh. In this set-up, there are no pure control districts. Rather, we compare the difference in the differences that exist between flood-prone and not and pre/post 2010 to establish a counterfactual. Our hypothesis is that while and . The main identifying assumption is that if STRVs had not been released, yields in flood-prone districts would have followed the same trend as yields in non-flood-prone districts. Obviously, flood-prone districts are likely to have lower yields than non-flood-prone districts. Thus we assume a mean shift in yields, as without STRVs mean yields are likely to be lower in flood-prone districts than in districts that are not prone to flooding, though the existence of parallel trends remains unobservable.
@@
As with the event study set-up, the DID model is not without its shortcomings. In particular, in the DID model we assume that farmers in all districts have access to STRVs starting in 2010. This assumption avoids endogeneity concerns about targeting of certain districts for seed but, based on the data, we know that STRVs were not universally available in 2010. In fact, in our administrative data, only one district had seed locally produced and distributed that early. And in the RMS data, adoption in 2014 was at eight percent. This suggests that, like any agricultural technology, roll out, dissemination, and adoption was gradual and took time. Th end results is that our use of a dummy for seed release in 2010, while exogenous, likely overestimates the number of districts βtreatedβ with STRV seed. A second shortcoming is that while a relative measure like flood-prone might drive adoption of STRVs and impact EVI, it is a coarse measure. It is also subjective and may not reflect agronomically important flood events.
@@
One main difference is that the TWFE approach uses administrative data on cumulative seed available in each district in each year instead of a binary indicator for if the observation comes before or after 2010. The other main difference is that we use EO data on actual flooding in a given year instead of a categorical variable for if the district is flood-prone or not. All other variables are the same as in the DID model.
@@
The benefit of the TWFE approach is that it uses available data on seed availability and flooding. In the event study, we controlled for flooding but our variable of interest measured changes in EVI in each year after seeds were introduce not changes to EVI when seeds were introduced and flooding occurred. This likely underestimates the effect of STRVs. In the DID model, we interact indicators for flooding and the introduction of STRVs but our variable of interest considers all districts as having equal access to seeds in every year since 2010. This also likely underestimates the effect of STRVs. Our TWFE model addresses these issues but it comes with the caveat that if seed availability is a function of time-varying, district-level events it is endogenous. Recall that we control for time invariant district characteristics, like soil quality and climate, as well as district invariant time events, like changes to national policy. So, the administrative seed data will only be endogenous if dissemination changed based on some time varying, district-level event, such as targeting a district for increased dissemination in the year after an unexpectedly severe flood.

### v2

@@ 3.3 Creation of District-Level EO Dataset
The creation of the district-level panel data follows the workflow summarized in FigureΒ 4. We start by using the rice area maps for Aman season for each year to mask the flood maps, so that we only consider inundation levels for rice area. We also use the rice maps to mask EVI data from MODIS so that we only consider EVI values for rice area.
@@
Given the resolution of the data, it is not possible to measure EVI on a rice field for use in predicting yields as in Lobell etΒ al. (2020). The rice maps exist at a spatial resolution of 250m while the mean size of a rice plot in Bangladesh is around two thirds of a hectare. Nor is it possible to measure the exact, agronomically relevant inundation on a given rice field since the flood maps exist at 500m resolution. To overcome this data limitation, we calculate a variety of EVI and flood metrics to ensure our proxies for relevant yield and flooding are robust. For EVI, we calculate the maximum value of EVI for the season, the mean and median values, and the cumulative EVI (sum of EVI throughout the season). FigureΒ 5 summarizes how each of these values change in a district before and after the introduction of STRVs into the district.
As flooding is the more sensitive metric (relative to adoption) based on the field trial data and our Monte Carlo simulations, we generate a large number of flood metrics. We create variation in the flood maps by 1) varying the amount of water above baseflow that we consider to be a flood and 2) varying the start and end day of the flood window (length) as well as the days within that window (duration) necessary to be considered a Goldilocks flood. To vary the amount of flooding, we draw the distribution of flooding across the entire time series. We then define a pixel as being flooded if its fractional value is in the top five percent of values ( of the time series is below that value) and zero otherwise. We repeat this process for every five percentile increment up to the percentile ( of the time series is below that value) to create 19 different flood quantiles.
Panel A of FigureΒ 6 shows the results for all of Bangladesh. Flooding at Q 5 ( percentile) represents a very low threshold for flooding while flooding at Q 95 ( percentile) represents extreme flooding, which is why the map at Q 95 is very blue - almost all the pixel has to be full (i.e., the fraction index is close to one) before it registers as a flood. Panel B aggregates the pixel-level fractional flood measure to the district-level and, for each quantile, calculates how many years that district experienced a flood in the largest possible flood window (5-20 days). For flooding at quantiles between Q 5 and Q 50, no district experiences a flood event in the flood window. This is because the threshold is so low districts look flooded for greater than 20 days. At high thresholds, like Q 95, where only extreme floods get counted, many districts experience numerous floods within the flood window. In fact, at Q 95 the coastal area and Northeast region appear to experience flood events in nearly every one of the 20 years of study. Panel C draws the distribution of the district-level days of consecutive flooding at each quantile. At low quantiles (Q 5), almost any amount of water above baseflow is considered a flood, so the length of consecutive days of flooding stretches for almost the entire rice season. At high quantiles (Q 95), inundation has to be extremely severe to register as a flood, meaning few of these occur and when they do they last for a very small consecutive number of days.
To vary the length of flooding, we create a set of potentially relevant flood windows that differ in start and end data and thus differ in length. The flood window can start at any day between five and 15 days after onset and can end at any day between 10 and 20 days after onset. We then bound the number of days that can be in the flood window between five and 15, so the largest flood window is the 15 day window starting on day five and ending on day 20. There are 11 equally small flood windows, each of five day length, ranging in start date from day five to day 15. So, for example, one flood window starts at day six and end at day 11 for a length of five days. Another starts at day five but ends at day 18 for a length of 13 days. And yet another starts at day 12 and ends at day 20 for a length of eight days. Systematically varying the window like this produces 66 potentially relevant flood windows.
We then intersect the number of days of consecutive flooding at each quantile with the flood window. This intersection produces 1,254 possible measures of the βjust rightβ amount of flooding. However, many of the intersections are empty, in that no district experiences flooding that falls within a given window. As the lower panel of FigureΒ 6 shows, at lower quantiles, where inundation needs to be minimal in order to be considered a flood, the duration of a flood lasts almost the entire season. Eliminating these empty intersections gives us 656 flood windows in which there is variation across space and time in the districts.
@@
The final piece of data that we integrate into our district-level panel is administrative data on seed dissemination. From IRRI and the Bangladesh Rice Research Institute (BRRI), we developed an exhaustive list of all seed production organizations, both private and public. While both IRRI and BRRI have been instrumental in developing submergence tolerant seeds varieties, neither is engaged in seed multiplication, which is conducted by a number of public and private entities. Investment in the seed system helped promote and disseminate STRVs from 2010 through 2019. We visited each seed producer/distributor and obtained disaggregate (district) level data on 1) the STRV variety names being multiplied or sold/distributed in the district, 2) the amount (tons) of each variety being multiplied in the district, and 3) the amount (tons) of each variety being sold/distributed in the district. FigureΒ 7 aggregates data from the 64 districts up to the division level and graphs the cumulative amount of seed available over time. We use the cumulative amount of seed available in each district in each year to account for the growing availability of STRVs and the fact that farmers recycle seed or obtain seed from outside their district.
@@ 3.4 Data Limitations
Before moving to our empirical method, it is important to note and discuss the data limitations in the study. In an ideal situation, researchers would have collected baseline data from rice growing households prior to the release of STRV seeds and would have followed-up with these households periodically over the years. Unfortunately, no such panel with pre-release baseline data exists. This motivates our use of EO data to reconstruct how the EVI signal responds to flooding pre-STRV release and then to examine how that signal changes post release.
@@
In terms of ideal EO data, one would like data from a sensor that is at both a high spatial and temporal resolution and that can penetrate clouds. Additionally, one would like to have ground truth data in terms of both flooding and rice area/yields across both space and time. As highlighted in the previous sections, such data does not exist, mainly because high resolution, cloud penetrating EO sensors, like Sentinel-1, did not exist in the public domain prior to the introduction of STRVs. Thus, there is no way to construct a baseline at both a high spatial and temporal resolution. Additionally, baseline ground truth data do not exist because no baseline survey was conducted. This poverty of ideal data motivates our approach to constructing rice area maps using Google Earth images and flood maps using our CNN-LSTM model.
@@
In developing our approach, we prioritized products with a high temporal resolution (MODIS) at the expense of products with a high spatial resolution (Landsat) because of the importance of capturing transitory flood events and EVI values immediately prior to harvest. We are aware of this trade-off and the limitation it creates: our EO dataset is constructed at 500m, which is an area equal to 25ha. In a cropping system like rice in Bangladesh, where the average rice area for a household is less than one hectare, this spatial resolution is not ideal. However, considering the types of measurement error introduced by the trade-off leads us to believe that prioritizing temporal over spatial resolution is the preferred approach. We have explored conducting the analysis using Landsat 30m data. However, Landsatβs temporal resolution is 16 days, meaning it frequently misses known flood events and EVI readings at harvest are frequently missed. Given this temporal resolution, it is possible to under count by missing a flood or harvest but it is not possible to over count. This means that the measurement error introduced is non-classical and while the direction of the bias is known, the size of the bias is unknown. By contrast, the relatively course spatial resolution of MODIS, combined with the ground truth data, and our machine learning classification approach means that we may misallocate a pixel to rice/not-rice and to flood/not-flood, but that misallocation is likely random. This means the measurement error introduced should be classical - increasing the variance or noise in the data but not introducing any systematic bias. Essentially, the spatial fuzziness of MODIS allows us to make guided inference on whether or not a pixel is rice/not-rice or flooded/not-flooded while the missing temporal images of Landsat are not amenable to making inference about what happened in those missing pixels. This spatial fuzziness can be measured in terms of accuracy metrics for our classification algorithms (RF and CNN-LSTM), which are in the range of -.
@@
Beyond the issue of coarse spatial resolution, we also lack objectively measured harvest information for training an ML algorithm on to predict yields. This forces us to rely on off-the-shelf EVI measures as proxies for yields, though this approach is not without its critics. A large body of research demonstrates strong correlation between EVI, and particularly maximum EVI, and yields (Mkhabela etΒ al., 2011; Liu etΒ al., 2020; Lobell etΒ al., 2020). However, the accuracy of using EVI as a proxy for yields varies substantially by location (Azzari etΒ al., 2017) and crop, with EVI performing particularly poorly for rice in the U.S. (Johnson, 2016). For rice in Bangladesh, the correlation between EO derived metrics and yield can vary from - (Islam etΒ al., 2021). However, Islam etΒ al. (2021) relied on government reported area yield statistics, a notoriously inaccurate source for measures of yield in low-income countries (Kilic etΒ al., 2021), where statistical departments are often under-resourced and politicized. We are unaware of tests of accuracy of EVI with rice yields in Bangladesh that rely on objective measures of yield.
@@
While these data limitations are not unique to our study, and would be present in any attempt to measure impact of an agricultural technology introduced over a decade ago, they should be kept in mind when interpreting our results.
4 Empirical Method
Our empirical method follows that laid out in Gollin etΒ al. (2021) to estimate the impact of Green Revolution technologies on yields. Our empirical strategy relies on the release dates of various varieties and the staggered accumulation of seed in each district, which we argue is plausibly exogenous to rice yields. We estimate the effects of STRVs on EVI as a proxy for yields using Aman season data at the district-level. We use a variety of econometric methods (event study, difference-in-difference, two-way fixed effects), each with its own strengths and weaknesses regarding identifying assumptions and the potential for attenuation bias due to mismeasurement.
We start by estimating a simple event study model:
@@ 4 Empirical Method
where indexes districts and indexes time. The two terms and denote district and time fixed effects, meaning that only within-district time variation in relative EVI remains. The district fixed effects control for all district-specific time-invariant variation and the time fixed effects control for all time-specific district-invariant variation. We also include a control for the percentage of a district that was flooded, which varies by district and time.
@@
There are two shortcomings to the event study approach. One is that STRVs reach different districts in different years. It is plausible that STRV seeds reached more flood-prone districts, or districts with more advanced farming techniques, earlier than other districts. If these characteristics are time-invariant, our district fixed effects will pick them up. However, one can imagine situations where time-variant, district-specific events accelerated the first introduction of STRV seed. While we explicitly control for contemporaneous flooding, flood events in the recent past might make the year a district got STRVs no longer exogenous. A second shortcoming is that STRVs are likely to increase EVI only very slightly over time, as a change in EVI would only occur when there was a Goldilocks flood and sufficient seed was available. If a district did not experience such a flood in a year after STRVs were released, or had very low quantities of seed, we would expect no change in EVI. Thus, the event study will underestimate the effects of STRV release date on EVI.
@@
To address these two issue, we next estimate a simple difference-in-difference (DID) model:
@@
where is an indicator equal to one in the years after 2010, they year the Bangladesh government approved STRVs, and zero otherwise. Similarly, equals one in districts that are prone to flooding and zero in districts that are not prone to flooding. To define a district as flood-prone or not, we first calculate the median amount of flooding across all districts and all years. We then calculate the median flood value in each district across the time series. We categorize a district as flood-prone if its median value is above the median value in all the data. The remaining terms are as defined in EquationΒ (4.1).101010Note that because we have converted both adoption and flood to binary indicators that do not vary over time, our DID estimator is not subject to the recently exposed potential biases of DID in complicated settings. For a summary of what settings complicate DID estimation and the numerous ways to correct for bias, see Roth etΒ al. (2023).
@@
We expect EVI in flood-prone districts to increase after 2010, which was when STRVs first became available in Bangladesh. In this set-up, there are no pure control districts. Rather, we compare the difference in the differences that exist between flood-prone/not and pre/post 2010 to establish a counterfactual. Our hypothesis is that while and . The main identifying assumption is that if STRVs had not been released, yields in flood-prone districts would have followed the same trend as yields in not-flood-prone districts. Obviously, flood-prone districts are likely to have lower yields than not-flood-prone districts. Thus we assume a mean shift in yields, as without STRVs mean yields are likely to be lower in flood-prone districts than in districts that are not prone to flooding, though the existence of parallel trends remains unobservable.
@@
As with the event study set-up, the DID model is not without its shortcomings. In particular, in the DID model we assume that farmers in all districts have access to STRVs starting in 2010. This assumption avoids endogeneity concerns about targeting of certain districts for seed but, as shown in FigureΒ 7, we know that STRVs were not universally available in 2010. In fact, in our administrative data, only one district had seed locally produced and distributed that early. This suggests that, like any agricultural technology, for STRVs the process of roll out, dissemination, and adoption was gradual and took time. The end result is that our use of a dummy for seed release in 2010, while exogenous, likely overestimates the number of districts βtreatedβ with STRV seed, attenuating the true treatment effect. A second shortcoming is that while a relative measure like flood-prone might drive adoption of STRVs and impact EVI, it is a coarse measure that does not vary over time.
@@
One main difference is that the TWFE approach uses administrative data on cumulative seed available in each district in each year instead of a binary indicator for if the observation comes after 2010. The other main difference is that we use measures of actual flooding in a given year instead of a categorical variable for if the district is flood-prone or not. All other variables are the same as in the DID model.111111The use of data on the staggered rollout of STRV seed means our TWFE model may be subject to the recently exposed potential biases of TWFE in complicated settings. However, as Wooldridge (2021) argues, the main issue with the older TWFE literature was that it did not account for heterogeneity that impacted both treatment effects and common trends. In our model we have explicitly accounted for heterogeneity by interacting treatment (STRV) with flooding in the belief that both treatment effect and common trends will vary by flooding. Thus, our models already accounts for the existence of heterogeneity in treatment and trends that many of the new methods are designed to address.
@@
The benefit of the TWFE approach is that it uses available data on seed distribution and flooding. In the event study, we controlled for flooding but our variable of interest measured changes in EVI in each year after seeds were introduce not changes to EVI when seeds were introduced and flooding occurred. This likely underestimates the effect of STRVs. In the DID model, we interact indicators for flooding and the introduction of STRVs but our variable of interest considers all districts as having equal access to seeds in every year since 2010. This also likely underestimates the effect of STRVs. Our TWFE model addresses these issues but it comes with the caveat that if seed availability is a function of time-varying, district-level events it may be endogenous. Recall that we control for time-invariant district characteristics, like soil quality and propensity to flood, as well as district-invariant time events, like changes to national agricultural policy. So, the administrative seed data will only be endogenous if dissemination changed based on some time-varying, district-level event, such as targeting a district for increased dissemination in the year after an unexpectedly severe flood.